OpenStudy (ebonyserrano):

Write the equation of the line that passes through (3, 4) and (2, –1) in slope-intercept form. A.y = 3x – 7 B. y = 3x – 5 C.y = 5x – 11 D. y = 5x – 9

563blackghost (563blackghost):

Slope formula: \(\huge{Slope=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}}\) Plug in... \(\huge{Slope=\frac{-1-4}{2-3}}\)

OpenStudy (ebonyserrano):

So would it be B?

563blackghost (563blackghost):

Not quite. We need to find slope once we do we plug it into the slope-intercept formula.... \(\huge{y=mx+b}\) \(\Large{m=slope}\) \(\Large{b=y-intercept}\) ~~~~~~~~~~~~~~~~~~~~~~~~~~ \(\huge{Slope=\frac{-5}{-1}=5}\) ~~~~~~ \(\Huge{y=5x+b}\) We would now find b so we would input a set of points into y and x....lets use (3,4)... \(\huge{4=5(3)+b}\)

563blackghost (563blackghost):

What would b equal?

OpenStudy (ebonyserrano):

y?

563blackghost (563blackghost):

To find b we would need to input a set of points into the open `y` and `x`....We will use `(3,4)` for this.... \(\huge{y=5x+b \rightarrow 4=5(3)+b}\) We have our original equation in which we apply the points...We now would simplify.... \(\huge{4=5(3)+b \rightarrow 4=15+b}\) We would then subtract `15` to the other side... \(\huge{(4-15=(15-15)+b \rightarrow~ ?}\) What would that equal?

OpenStudy (ebonyserrano):

11

OpenStudy (ebonyserrano):

?

563blackghost (563blackghost):

Correct :D Now we would input 11 as b and get our final equation... \(\Huge\color{red}{y=5x-11}\)

OpenStudy (ebonyserrano):

thank you!

563blackghost (563blackghost):

np :D